Solve by Factoring:
x^2-9x=-20
x^2-9x+20=0
x^2-4x-5x+20=0
x(x-4)+(-5)(x-4)=0
(x-5)(x-4)=0
(x-5) or (x-4)
x=5 or x=4
Is this Correct? Thank you

Answer by mukhopadhyay(490)

It is mostly corect
x^2-9x=-20
x^2-9x+20=0
x^2-4x-5x+20=0
x(x-4)+(-5)(x-4)=0
(x-5)(x-4)=0
(x-5)=0 or (x-4)=0............(This is where you made a small mistake)
x=5 or x=4